OpenStudy (anonymous):

Solve the system of equations by using the substitution method. 18x-6y=4 x+y=0 I tried substituting the second equation into the first one, but my answer was wrong. Please help. Thanks!

OpenStudy (anonymous):

what variable did you substitute? what equation did you use to substitute? ( yes i know you used the 2nd equation but you can just use it as is) any way you can show your work?

OpenStudy (anonymous):

cant*

OpenStudy (anonymous):

I put y=x+0 in the first equation, then multiplied as normal and I came out with a decimal.

OpenStudy (anonymous):

ok, thats where you made your mistake if we take the 2nd equation x+y=0 and solve for "y" that would mean we would have to subtract x from both sides x+y = 0 x+y - x = 0 -x and we would get y= -x instead of y=x

OpenStudy (anonymous):